# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ 7 Discussion
Finally, as we were finalizing the manuscript for submission during the summer of 2024, the incumbent President Joe Biden has dropped out of the 2024 U.S. presidential election in late July of 2024; our original analysis plan assumed that President Biden is the Democratic Party’s nominee for the presidency. While we believe the interpretations from our analysis about ads will still be plausible since Trump is the nominee for the Republication party and the digital ad campaign consists of negative ads against Trump, we caution readers from over-interpreting the results. Notably, our calibration procedure based on blue-collar and non blue-collar states could under-estimate the dramatic shift in electoral context after Biden dropped out of the race and the consequences of this unprecedented event in American politics.
@@
This Appendix provides supplementary materials for the manuscript titled “Transfer Learning Between U.S. Presidential Elections: How should we learn from a 2020 ad campaign to inform 2024 campaigns?”. Section A details the alternative formulations and extensions of the sensitivity model (3.1). Sections B and C provide additional details and proofs for inference procedures based on the OR estimator and the EIF-based estimator introduced in Sections 4.1 and 4.2 of the main text, respectively. Section D provides details of the calibration procedure of sensitivity parameters. Section E provides additional results for analyzing the ad effect in PA in 2024. Section F repeats the analysis on PA when excluding voters in PA, WI, MI from the source data (i.e., restricting the source data to voters from NC and AZ). Section G presents a simulation study where the simulation data were generated to mimic the data from Aggarwal et al., (2023).
@@ A.1 Exponential Tilting for Continuous Outcomes
The proposed sensitivity model (3.1) is not limited to binary outcomes. It can be equivalently expressed for a general, possibly continuous outcome with support ${\cal Y}$. Suppose the conditional density of the potential outcome on the target population is shifted from that of the source by an exponential tilting shift,
@@
where $\propto$ represents “proportional to” and $p_{Y^{(a)}\mid{\bf V},S=s}$ represents the conditional probability density function of $Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=s$ for $s=0,1$. When $\gamma_{a}=0$, (A.1) reduces to $p_{Y^{(a)}\mid{\bf V},S=0}(y_{a}\mid{\bf v},S_{i}=0)=p_{Y^{(a)}\mid{\bf V},S=0}(y_{a}\mid{\bf V}_{i}={\bf v},S_{i}=1)$ and thereby transportability (Assumption 2.3) holds. When $\gamma_{a}\neq 0$, $\gamma_{a}$ measures the violation to the transportability assumption by the degree in shifts of the conditional densities.
@@ Appendix B Details and Proofs for the Outcome Regression Based Estimation
B.1 Details for the Bootstrap
@@
3:Step 2: Estimate $\widehat{\theta}_{\textnormal{OR,a}}(\gamma_{a})$ as in (4.1).
@@
$\widehat{Q}^{*}(\tau)=\inf_{t}\left\{\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}(\widehat{\theta}_{\textnormal{OR}}^{*,b}(\gamma_{a})\leq t)\geq\tau\right\},\forall\tau\in(0,1).$
@@
B.2 Regularity Conditions for the Bootstrap
@@
$\phi_{a}({\bf V}_{i},\theta_{a}(\gamma_{a}),\boldsymbol{\eta}_{a})$ $=\dfrac{\exp(\gamma_{a})\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})}{\exp(\gamma_{a})\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})+1-\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})}-\theta_{a}(\gamma_{a}).$
@@
Suppose Assumptions 2.1 and 2.2 hold and $n_{s}\asymp n_{t}$. Also suppose $\rho_{a}({\bf v},\boldsymbol{\eta}_{a})$ is twice differentiable with respect to $\boldsymbol{\eta}_{a}$ and $\widehat{\boldsymbol{\eta}}_{a}$ is an asymptotically linear estimate of $\boldsymbol{\eta}_{a}$ with some influence function ${\bf g}_{a}({\bf O},\boldsymbol{\eta}_{a})$; i.e., $\sqrt{n}(\widehat{\boldsymbol{\eta}}_{a}-\boldsymbol{\eta}_{a})=\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}{\bf g}({\bf O}_{i},\boldsymbol{\eta}_{a})+o_{p}(1)$. If $\theta_{a}(\gamma_{a})\in\Theta$ where $\Theta$ is open and compact, then $\widehat{\theta}_{\textnormal{OR,a}}(\gamma_{a})\to_{p}\theta_{a}(\gamma_{a})$ and $\widehat{\theta}_{\textnormal{OR,a}}(\gamma_{a})$ is asymptotically linear with influence function
@@
$\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})}{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}-\theta_{1}\right\rVert=o_{p}(1).$
@@
$\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})}{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\theta_{1}\right\rVert$
$\leq$ $\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})}{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})}{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right\rVert$
$+\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})}{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}-\theta_{1}\right\rVert$
$\leq$ $\dfrac{\exp(\gamma_{1})}{\min\{1,\exp(\gamma_{1})\}}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\widehat{\rho}_{1}({\bf V}_{i})-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\rho_{1}({\bf V}_{i})\right\rVert+o_{p}(1),$
@@ Appendix C Details and Proofs for the EIF-Based Estimation
C.1 Estimating the Density Ratio
By Bayes rule, $w({\bf v})$ can be expressed in terms of the selection probability $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ for ${\bf v}\in{\cal V}$,
$w({\bf v})=\dfrac{\mathbb{P}(S_{i}=1)}{\mathbb{P}(S_{i}=0)}\dfrac{\mathbb{P}(S_{i}=0\mid{\bf V}_{i}={\bf v})}{\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})}.$

### v2

@@ 7 Discussion and future work
Finally, as we were finalizing the manuscript during the summer of 2024, the incumbent President Joe Biden has dropped out of the 2024 U.S. presidential election in late July of 2024; our original analysis plan assumed that President Biden is the Democratic Party’s nominee for the presidency. While we believe the interpretations from our analysis is still plausible since Trump was the nominee for the Republication party and the digital ad campaign consisted of negative ads against Trump, we caution readers from over-interpreting the results. Notably, our calibration procedure based on the rust belt and the sun belt states could under-estimate the dramatic shift in electoral context after Biden dropped out of the race and the consequences of this unprecedented event in American politics.
@@ A.1 Exponential Tilting for Continuous Outcomes
The proposed sensitivity model (2) is not limited to binary outcomes. It can be equivalently expressed for a general, possibly continuous outcome with support ${\cal Y}$. To ease communication, we let $\gamma_{a}={\textnormal{log}}(\Gamma_{a})\in\mathbb{R}$. Suppose the conditional density of the potential outcome on the target population is shifted from that of the source by an exponential tilting shift,
@@
where $\propto$ represents “proportional to” and $p_{Y^{(a)}\mid{\bf V},S=s}$ represents the conditional probability density function of $Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=s$ for $s=0,1$. When $\gamma_{a}=0$ (i.e., $\Gamma_{a}=1$), (A.1) reduces to $p_{Y^{(a)}\mid{\bf V},S=0}(y_{a}\mid{\bf v},S_{i}=0)=p_{Y^{(a)}\mid{\bf V},S=0}(y_{a}\mid{\bf V}_{i}={\bf v},S_{i}=1)$ and thereby transportability (Assumption 2.3) holds. When $\gamma_{a}\neq 0$, $\gamma_{a}$ measures the violation to the transportability assumption by the degree in shifts of the conditional densities.
@@ Appendix B Details and Proofs for the Outcome Regression Based Estimation
B.1 Estimation of $\rho_{a}$
We verify that when ${\cal X}$ and ${\cal V}$ are discrete and $\pi$ and $\mu_{a}$ are estimated by group averages, then estimators motivated from (8), (9) and (10) are equivalent. To be explicit, for a given ${\bf x}\in{\cal X}$, the estimates of $\pi$ and $\mu_{a}$ are
$\widehat{\pi}({\bf x})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}A_{i}\mathds{1}({\bf X}_{i}={\bf x})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf X}_{i}={\bf x})},$
$\widehat{\mu}_{a}({\bf x})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})Y_{i}}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})},$
respectively. The equalities in (8), (9), (10) suggest an outcome regression typed estimator $\widehat{\rho}_{a}^{\text{OR}}$, an inverse probability weighting estimator $\widehat{\rho}_{a}^{\text{IPW}}$, and The equality in (8) suggests an outcome regression typed estimator that we denote as an augmented inverse probability weighting estimator $\widehat{\rho}_{a}^{\text{AIPW}}$, respectively, where for ${\bf v}\in{\cal V}$,
$\widehat{\rho}_{a}^{\textnormal{OR}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},$
$\widehat{\rho}_{a}^{\textnormal{IPW}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\left\{\dfrac{A_{i}Y_{i}}{\widehat{\pi}({\bf X}_{i})}+\dfrac{(1-A_{i})Y_{i}}{1-\widehat{\pi}({\bf X}_{i})}\right\}\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},$
$\widehat{\rho}_{a}^{\textnormal{AIPW}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}({\bf X}_{i})}+\dfrac{(1-A_{i})}{1-\widehat{\pi}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}({\bf X}_{i})\}+\widehat{\mu}_{a}({\bf X}_{i})\right]\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}.$
Lemma B.1
When ${\cal X}$ and ${\cal V}$ are discrete, $\widehat{\rho}_{a}^{\textnormal{OR}}({\bf v})=\widehat{\rho}_{a}^{\textnormal{IPW}}({\bf v})=\widehat{\rho}_{a}^{\textnormal{AIPW}}({\bf v})$ for any ${\bf v}\in{\cal V}$.
Proof of Lemma B.1. Without loss of generality, we prove for the case when $a=1$.
First, we show that $\widehat{\rho}_{1}^{\textnormal{OR}}({\bf v})=\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$. We can simplify $\widehat{\rho}_{1}^{\textnormal{OR}}({\bf v})$ as
$\widehat{\rho}_{1}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}$
$=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\cdot\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})Y_{j}}{\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{k}=1,{\bf X}_{k}={\bf X}_{i})}}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})Y_{j}}{\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{k}=1,{\bf X}_{k}={\bf X}_{i})}\right\}.$ (B.1)
We can simplify $\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$ as
$\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\mathds{1}(A_{i}=1)Y_{i}}{\widehat{\pi}({\bf X}_{i})}\right\}\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\cdot\sum_{i\in{\cal I}_{s}}\left[\dfrac{\mathds{1}(A_{i}=1)Y_{i}}{\widehat{\pi}({\bf X}_{i})}\right]$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\cdot\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\mathds{1}(A_{i}=1)Y_{i}}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1)\mathds{1}({\bf X}_{j}={\bf X}_{i})/\left\{\sum_{k\in{\cal I}_{s}}\mathds{1}({\bf X}_{k}={\bf X}_{i})\right\}}\right\}$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\cdot\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{i}=1)Y_{i}\mathds{1}({\bf X}_{k}={\bf X}_{i})}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1)\mathds{1}({\bf X}_{j}={\bf X}_{i})}\right\}$ (B.2)
[…]
Next, we show that $\widehat{\rho}_{1}^{\textnormal{AIPW}}({\bf v})=\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$.
$\widehat{\rho}_{1}^{\textnormal{AIPW}}({\bf v})-\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})=$ $\dfrac{\sum_{i\in{\cal I}_{s}}\left\{-\dfrac{A_{i}}{\widehat{\pi}({\bf X}_{i})}+1\right\}\widehat{\mu}_{a}({\bf X}_{i})\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},$
where the numerator is
$\sum_{i\in{\cal I}_{s}}\left\{-\dfrac{\mathds{1}(A_{i}=1)}{\widehat{\pi}({\bf X}_{i})}+1\right\}\widehat{\mu}_{a}({\bf X}_{i})\mathds{1}({\bf V}_{i}={\bf v})$
$=$ $\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})\left\{\dfrac{-\mathds{1}(A_{i}=1)+\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{k\in{\cal I}_{s}}}\mathds{1}({\bf X}_{k}={\bf X}_{i})}{\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{k\in{\cal I}_{s}}\mathds{1}({\bf X}_{k}={\bf X}_{i})}}\right\}$
$=$ $\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})\dfrac{-\mathds{1}(A_{i}=1)\sum_{k\in{\cal I}_{s}}\mathds{1}({\bf X}_{k}={\bf X}_{i})+\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}$
$=$ $\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})\dfrac{-\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{i}=1)\mathds{1}({\bf X}_{k}={\bf X}_{i})+\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}$
$=$ $0$
Therefore, $\widehat{\rho}_{1}^{\textnormal{AIPW}}({\bf v})=\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$.
$\square$
B.2 Details for the Bootstrap
@@
3:Step 2: Estimate $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$ as in (4).
@@
$\widehat{Q}^{*}(\tau)=\inf_{t}\left\{\frac{1}{B}\sum_{b=1}^{B}\mathds{1}(\widehat{\theta}_{\textnormal{OR}}^{*,b}(\Gamma_{a})\leq t)\geq\tau\right\},\forall\tau\in(0,1).$
@@
B.3 Regularity Conditions for the Bootstrap
@@
$\phi_{a}({\bf V}_{i},\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})$ $=\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})}{\Gamma_{a}\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})+1-\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})}-\theta_{a}(\Gamma_{a}).$
@@
Suppose Assumptions 2 and 3 hold and $n_{s}\asymp n_{t}$. Also suppose $\rho_{a}({\bf v},\boldsymbol{\eta}_{a})$ is twice differentiable with respect to $\boldsymbol{\eta}_{a}$ and $\widehat{\boldsymbol{\eta}}_{a}$ is an asymptotically linear estimate of $\boldsymbol{\eta}_{a}$ with some influence function ${\bf g}_{a}({\bf O},\boldsymbol{\eta}_{a})$; i.e., $\sqrt{n}(\widehat{\boldsymbol{\eta}}_{a}-\boldsymbol{\eta}_{a})=\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}{\bf g}({\bf O}_{i},\boldsymbol{\eta}_{a})+o_{p}(1)$. If $\theta_{a}(\Gamma_{a})\in\Theta$ where $\Theta$ is open and compact, then $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})\to_{p}\theta_{a}(\Gamma_{a})$ and $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$ is asymptotically linear with influence function
@@
$\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}-\theta_{1}\right\rVert=o_{p}(1).$
@@
$\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})}{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\theta_{1}\right\rVert$
$\leq$ $\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})}{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right\rVert$
$+\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}-\theta_{1}\right\rVert$
$\leq$ $\dfrac{\Gamma_{1}}{\min\{1,\Gamma_{1}\}}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\widehat{\rho}_{1}({\bf V}_{i})-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\rho_{1}({\bf V}_{i})\right\rVert+o_{p}(1),$
@@ Appendix C Details and Proofs for the EIF-Based Estimation
In this section we provide details and proofs for the EIF-based estimator $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ proposed in Section 4.2 of the main text.
C.1 Implementation Details
For the EIF-based estimator $\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$, we remark that the second sum may be replaced by a counterpart that does not use sample splitting, i.e., $\dfrac{1}{|{\cal I}_{t}|}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})}$, and the resulting estimator will have the same asymptotic distribution as $\widehat{\theta}^{(k)}_{\textnormal{EIF,a}}$. For simplicity of presentation and implementation, we focus our attention on $\widehat{\theta}^{(k)}_{\textnormal{EIF,a}}$ throughout as the first sum in the EIF estimator requires sample splitting.
A step-by-step implementation of the EIF-based estimation is provided in Algorithm 2.
Algorithm 2 EIF-Based Estimation with Cross-Fitting
1:${\cal I}_{s}$, ${\cal I}_{t}$; integer $K\geq 2$; sensitivity parameters $\Gamma_{1},\Gamma_{0}$; confidence level $(1-\alpha)$.
2:Step 1 (Partitioning): Randomly split ${\cal I}_{s}$ and ${\cal I}_{t}$ to ${\cal I}_{s,k}$ and ${\cal I}_{t,k}$, respectively, $1\leq k\leq K$.
3:Step 2 (Cross Fitting):
4:for $k$ in $1,2,\cdots K$ do
5:     With $\{{\cal I}_{s}\backslash{\cal I}_{s,k}\}\cup\{{\cal I}_{t}\backslash{\cal I}_{t,k}\}$, obtain $\widehat{\pi}^{(k)}({\bf X}_{i})$, $\widehat{w}^{(k)}({\bf V}_{i})$, $\widehat{\mu}^{(k)}_{a}({\bf X}_{i})$ and $\widehat{\rho}^{(k)}({\bf V}_{i})$.
6:     With ${\cal I}_{s,k}\cup{\cal I}_{t,k}$, calculate $\widehat{\theta}_{\textnormal{EIF}}^{(k)}(\Gamma_{0},\Gamma_{1})=\widehat{\theta}_{\textnormal{EIF,1}}^{(k)}(\Gamma_{1})-\widehat{\theta}_{\textnormal{EIF,0}}^{(k)}(\Gamma_{0})$ with $\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$ in Section 4.2.
7:end for
8:Step 3 (Building Estimator): Construct the estimator $\widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})=\dfrac{1}{K}\sum_{k=1}^{K}\widehat{\theta}_{\textnormal{EIF}}^{(k)}(\Gamma_{0},\Gamma_{1})$.
9:Step 4 (Variance Estimation): Construct the variance estimator
$\widehat{\sigma}^{2}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})=\dfrac{1}{K}\sum_{k=1}^{K}\left[\dfrac{1}{|{\cal I}_{k}|}\sum_{i\in{\cal I}_{k}}\left\{\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,1}}(\Gamma_{1}))-\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,0}}(\Gamma_{0}))\right\}^{2}\right].$
10:The EIF-based estimator $\widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})$ with a $(1-\alpha)$ confidence interval
$\left(\widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})-z_{\alpha/2}\widehat{\sigma}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})/\sqrt{n},\ \widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})+z_{1-\alpha/2}\widehat{\sigma}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})/\sqrt{n}\right),$
where $z_{\beta}$ is the $\beta$ quantile of the standard Normal distribution for any $\beta\in(0,1)$ and $\widehat{\sigma}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})=\sqrt{\widehat{\sigma}^{2}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})}$.
C.2 Estimating the Density Ratio
The estimation of the density ratio can proceed in two methods falling into two categories. The first category is to recognize the relationship between $w({\bf v})$ and $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ via the Bayes rule, i.e.,
$w({\bf v})$ $=\dfrac{\mathbb{P}(S_{i}=1)}{\mathbb{P}(S_{i}=0)}\dfrac{\mathbb{P}(S_{i}=0\mid{\bf V}_{i}={\bf v})}{\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})},$ (C.1)